OpenStudy (anonymous):

The width w of a rectangular swimming pool is 7x2. The area A of the pool is 7x3 – 42x2. What is an expression for the length of the pool?

OpenStudy (anonymous):

(7x^3 – 42x^2) /(7x^2) ?

OpenStudy (anonymous):

Area(A) = 7x^3 – 42x^2 = 7x^2(x-6)

OpenStudy (anonymous):

width(w) = 7x^2

terenzreignz (terenzreignz):

Medal, if you justify that, Corey :)

OpenStudy (anonymous):

Area = width*length Area/width = length or [7x^2(x-6)]/[7x^2] = length (x-6) = length x-6 = length (this is what they're looking for)

terenzreignz (terenzreignz):

Totally worth it :)

OpenStudy (anonymous):

NP (:

OpenStudy (anonymous):

@AngelCriner Am i Right ?

OpenStudy (anonymous):

Lol correct, thank you

OpenStudy (anonymous):

Welcome